OpenStudy (anonymous):

A class has 26 students. Each student has to present a project to the class. To decide the order, the teacher writes every letter in the alphabet on pieces of paper and puts it in a hat. What is the probability that the first letter drawn is not a vowel? A class has 26 students. Each student has to present a project to the class. To decide the order, the teacher writes every letter in the alphabet on pieces of paper and puts it in a hat. What is the probability that the first letter drawn is not a vowel?

OpenStudy (anonymous):

5/26

OpenStudy (anonymous):

5 vowels, and 26 letters all together.

OpenStudy (anonymous):

:O

OpenStudy (bibby):

NOT a vowel is 26-5/26

OpenStudy (anonymous):

1-5/26

OpenStudy (anonymous):

\[1-\frac{5}{26}=\frac{26}{26}-\frac{5}{26}\]

OpenStudy (bibby):

\[\huge \frac{ 26 }{ 26 } - \frac{ 5 }{ 26 }\] where 5/26 is the chance of getting a vowel

OpenStudy (anonymous):

the answer choices r ....A.\[\frac{ 9 }{ 26}\] B.\[\frac{ 5 }{ 26 }\] C.\[\frac{ 7 }{ 26 }\] D.\[\frac{ 21 }{ 26 }\]

OpenStudy (anonymous):

The answer is D.